OpenStudy (anonymous):

How do you do this? ? --> 43 Ca + 0 20 -1 e

sam (.sam.):

\[\huge _{q}^{p}X \rightarrow _{20}^{43}Ca+_{-1}^{0}e\] You know that mass and energy is conserved in a nuclear reaction, so equate them together p=43+0 q=20-1 ------------------------- p=43 q=19 \[\huge _{19}^{43}K\]